OpenStudy (anonymous):

x ^ 1/8 times y ^ 1/6 times z ^ 1/3 Use rational exponest to write as a single radical expression

jhonyy9 (jhonyy9):

\[\sqrt[8]{x}*\sqrt[6]{y}*\sqrt[3]{z}\]

jhonyy9 (jhonyy9):

\[\sqrt[8]{xy ^{2}z ^{5}}\]